A gas is initially at volume 40.0 L, pressure 600 mmHg and temperature 35 °C. Calculate the final pressure of the gas in atm if the volume and temperature are changed to 2500 mL and 225 K, respectively.

1
Expert's answer
2021-11-11T08:02:01-0500

Initial volume (V1) = 40 L ; pressure (P1) = 600 mmHg = 0.79 atm ;

Temperature (T1) = 35 + 273 K = 308 K

Final Volume (V2) = 2.5 L ; Temperature (T2) = 225 K ; Pressure of gas

We know that P1V1"\/" T1 = P2V2"\/" T2

Substitute all values above equation

(0.79) (40) "\/" (308) = (P2) (2.5) "\/" 225

Pressure of gas (P2) = 9.23 atm